What Is MIT SCG?

The MIT Student Cable Group is a growing organization of students producing a wide range of television programming and media production for the MIT community. Students play every role in television production: camera work, directing, writing, acting, and even repair work. Our programming includes movies, comedy, music videos and other musical presentations, and ideas that new people bring to us every week.

We’re best known for bringing web-based video-on-demand (called “the Button (tm)”) to MIT students. And, of course, we make the only programs on television about MIT. We provide equipment, advice, and help to members of the MIT community interested in creating their own programs. We are your ultimate channel.

Our driving forces are to:

1) teach, learn, and share information about

the technical aspects of television production

2) practice the art of television production

3) have fun

Unlike other college television stations which are controlled and staffed by communications schools, we are entirely student-operated. We are also the youngest major medium on campus.

Through the MIT Cable Television system’s almost 4,000 cable outlets, we can broadcast to and from all campus buildings and living groups. We broadcast on channels 35 and 36 over closed circuit cable.

We are always looking for new people. Even if you’ve never used a VCR before, you’re qualified, and welcome, to join the MIT Student Cable Group.

Our Programs

Pretty Late

Do you ever get the urge to talk about anything you want? Perhaps you would enjoy expressing your opinions across the entire MIT campus? Well, that’s what we do! Join us as our lively and informative discussions both entertain and offend you! There is no shortage of touchy-feely subjects, believe you me! Pretty Late is always current, and uniquely MIT’s. email: pretty-late@mit.edu

This Old ...

A short five minute piece every week focusing on the comedy of learning an everyday thing. It will be instructional to the mit college student yet allow the host to rant. First episode will probably be about boiling an egg. email: this-old@mit.edu

MIT2K

Putting our unique spin on movie commentating.

Contests

In past years, we have presented the 2.007 and 6.270 contests live from 26–100. These were our biggest and most popular projects every year until the organizers decided to pay (someone else) to tape it.

Other Projects

We televise concerts, public affairs programs, and live and taped events of interest. Watch out for Burly Bear network and CCTV.

Location of the Studio

How To Find Us

In Person

We will hold meetings this fall on the first Saturday of each month at 2pm in Room 9 –034. A schedule of taping sessions and other meetings is posted outside our office, Room 9 – 034, and new members are welcome to attend. Room 9 –034 is in the upper basement of building 9, which is locked after 5pm and on weekends. It may be entered through buildings 7, 13, or 35.

By Telephone

The station phone is (617) 252-1694 and can be reached on campus by dialing x2-1694. If no one is there, it will forward to the phone of one of the officers.

By Fax

We don’t have a fax, send us email.

By Interdepartmental Mail

Send interdepartmental mail to “MIT Student Cable Group, Room 9–034.”

By U. S. Mail

Address mail to MIT Student Cable Group, MIT Room 9–034, 77 Massachusetts Avenue, Cambridge, MA 02139. Large packages should be addressed to MIT Student Cable Group, Room 9–034, Building 3 Shipping Room, Cambridge, MA 02139.

Who To Talk To

Rachael Leventhal

General Mgr., President

lucky11@mit.edu

Philip Tracadas

Technical Mgr., Treasurer

tracadas@mit.edu

Greg Dancer

Production Manager

John Evans Josh Glazer Kwasi Kapori Joel Rosenberg

Producers

Randy Winchester

Group advisor, CAES

(617) 253-7431 Rm. 9–050

TV Studio

(617)252-1694, x2-1694

Rm. 9–034

http://mitv.mit.edu/

Officer list: student-cable@mit.edu Members list: mitv-all@mit.edu

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1999,

MIT Student Cable Group.

MIT Student ; Cable Group ; Channel 36